OpenStudy (anonymous):

A boat traveled downstream a distance of 80 mi and then came right back. If the speed of the current was 8 mph and the total trip took 5 hours and 20 minutes, find the average speed of the boat relative to the water.

OpenStudy (anonymous):

t = d/r Let c be the speed of the current and b be the speed of the boat. Solve the following for b:\[\frac{80}{b+8}+\frac{80}{b-8}=\frac{320}{60} \]b = 32 mph